# Equal contribution    Corresponding author

### v3

@@ Abstract
Text-to-Image (T2I) diffusion models have achieved remarkable success in image generation. Despite their progress, challenges remain in both prompt-following ability, image quality and lack of high-quality datasets, which are essential for refining these models. As acquiring labeled data is costly, we introduce AGFSync, a framework that enhances T2I diffusion models through Direct Preference Optimization (DPO) in a fully AI-driven approach. AGFSync utilizes Vision-Language Models (VLM) to assess image quality across style, coherence, and aesthetics, generating feedback data within an AI-driven loop. By applying AGFSync to leading T2I models such as SD v1.4, v1.5, and SDXL, our extensive experiments on the TIFA dataset demonstrate notable improvements in VQA scores, aesthetic evaluations, and performance on the HPSv2 benchmark, consistently outperforming the base models. AGFSync’s method of refining T2I diffusion models paves the way for scalable alignment techniques.
Keywords:
Text-to-Image Generation, Diffusion Models, Direct Preference Optimization, AI Feedback
Figure 1: We introduce AGFSync: a model-agnostic training algorithm that improves text-to-image (T2I) generation models’ faithfulness and coherence to text inputs and image aesthetics without human interventions. The images showcase a comparison of the results before and after fine-tuning SDXL with AGFSync.
@@ 1 Introduction
The advent of Text-to-Image (T2I) generation technology represents a significant advancement in generative AI. Recent breakthroughs have predominantly utilized diffusion models to generate images from textual prompts [23, 2, 17, 31]. However, achieving high fidelity and aesthetics in generated images continues to pose challenges, including deviations from prompts and inadequate image quality [30]. Addressing these challenges requires enhancing diffusion models’ ability to accurately interpret detailed prompts (prompt-following ability [2]) and improve the generative quality across style, coherence, and aesthetics.
@@
Efforts to overcome these challenges span dataset, model, and training levels. High-quality text-image pair datasets, as proposed in the data-centric AI philosophy, can significantly improve performance [32]. Therefore a high-quality image caption, and its corresponding image pair dataset is crucial in training [2]. At the model architecture level, advancements include the optimization of cross-attention mechanisms to improve model compliance [7]. These efforts, both at the dataset and model architecture levels, follow the traditional training paradigm of using elaborately designed models with specific datasets. In contrast, in the training domain, strategies inspired by the success of large language models, such as OpenAI’s ChatGPT [14], include supervised fine-tuning (SFT) and alignment stages. With a pretrained T2I diffusion model, enhancing the model for better image quality can be approached in either the SFT stage or the alignment stage. The former approach, as seen in the latest work DreamSync [26], fine-tunes the diffusion model through a selected image selection procedure where a Vision-Language Model (VLM) evaluates and then selects high-quality text-image pairings for further fine-tuning. However, DreamSync exhibits a lower prompt generation conversion rate and is limited by the intrinsic capabilities of the diffusion model, leading to uncontrollable data distribution in the fine-tuning dataset. The latter approach, DPOK [6], DDPO [3], and DPO [20] use reinforcement learning for alignment, while Diffusion-DPO [27] applies Direct Preference Optimization (DPO) for model alignment, modifying the original DPO algorithm to directly optimize diffusion models based on preference data. Yet, it only focuses on evaluating image quality from one aspect. Furthermore, existing methods mostly depend on extensive, quality-controlled labeled data.
@@
Addressing these requires a cost-effective, low-labor approach that minimizes the need for human-labeled data while considering multiple quality aspects of images. Leveraging AI in generating datasets and evaluating image quality can fulfill these gaps without human intervention. Through generating diverse textual prompts, assessing generated images, and constructing a comprehensive preference dataset, AGFSync epitomizes the full spectrum of AI-driven innovation—ushering in an era of enhanced data utility, accessibility, scalability, and process automation while simultaneously mitigating the costs and limitations associated with manual data labeling.
@@
More specifically, AGFSync aligns text-to-image diffusion models via DPO, with multi-aspect AI feedback generated data. The process begins with the preference candidate set generation, where LLM generates descriptions of diverse styles and categories, serving as high-quality textual prompts. Candidate images are then generated using these AI-generated prompt, therefore constructing candidate prompt-image pairs. Image evaluation and VQA data construction follow, using LLM to generate questions related to the composition elements, style, etc., based on its initial prompts. VQA scoring is conducted by inputting these questions into the VQA model to assess whether the diffusion model-generated images aesthetically follow the prompts, calculating accuracy as the VQA score. With combined weighted scores of VQA, CLIP, and aesthetics filtering, the preference pair dataset is established within the best and worst images. Finally, DPO alignment is applied to the diffusion model using the constructed preference pair dataset. The entire process leverages the robust capabilities of VLMs without any human engagement, ensuring a human-free, cost-effective workflow.
@@
    Our proposed framework AGFSync, aided by multiple evaluation scores, leveraging DPO fine-tuning approach, introduces a fully automated, AI-driven approach, which elevates fidelity and aesthetic quality across varied scenarios without human annotations.
@@
    Extensive experiments demonstrate that AGFSync significantly and consistently improves upon existing diffusion models in terms of adherence to text prompts and overall image quality, establishing the efficacy and transformative potential of our AI-driven data generation, evaluation and fine-tuning framework.
@@ 2.1 Aligning Diffusion Models Methods
Dataset-Level Approaches: A pivotal aspect of improving image generation models involves curating and fine-tuning datasets that are deemed visually appealing. Works by [22, 17] utilize datasets rated highly by aesthetics classifiers [24] to bias models towards generating visually appealing images. Similarly, Emu [5] enhances both visual appeal and text alignment through fine-tuning on a curated dataset of high-quality photographs with detailed captions. Efforts to re-caption web-scraped image datasets for better text fidelity are evident in [2, 25]. Moreover, DreamSync [26] improves T2I synthesis with feedback from vision-language image understanding models, aligning images with textual input and the aesthetic quality of the generated images.
Model-Level Enhancements: At the model level, enhancing the architecture with additional components like attention modules [7] offers a training-free solution to enhance model compliance with desired outputs. StructureDiffusion [7] and SynGen [21] also work on training-free methods that focus on model’s inference time adjustments.
Training-Level Strategies: The integration of supervised fine-tuning (SFT) with advanced alignment stages, such as reinforcement learning approaches like DPOK [6], DDPO [3], and DPO [20], demonstrates significant potential in aligning image quality with human preferences. Among these, Diffusion-DPO emerges as an RL-free method, distinct from other RL-based alignment strategies, effectively enhancing human appeal while ensuring distributional integrity [27].
Common drawbacks of them is the expensive fine-tuning dataset where most of these approaches rely on human-annotated dataset and human evaluation. This paradigm cannot support training an extensive and scalable diffusion model.
@@ 2.2 Image Quality Evaluation Methods
Evaluating image quality in a comprehensive manner is pivotal, integrating both automated benchmarks and human assessments to ensure fidelity and aesthetic appeal. The introduction of TIFA [9] utilize Visual Question Answering (VQA) models to measure the faithfulness of generated images to text prompts, setting a foundation for subsequent innovations. The CLIP score [8] builds upon CLIP [19] enables a reference-free evaluation of image-caption compatibility through the computation of cosine similarity between image and text embeddings, showcasing high correlation with human judgments without needing reference captions. PickScore [11] leverages user preferences to predict the appeal of generated images, combining CLIP model elements with InstructGPT’s reward model objectives [15] for a nuanced understanding of user satisfaction. Alongside, the aesthetic score [10] assesses images based on aesthetics learned from image-comment pairs, providing a richer evaluation that includes composition, color, and style.
@@ 3 Methodology
Figure 2: Overall pipeline of AGFSync, which mainly encompasses 3 steps: (1) Preference Candidate Set Generation, (2) Preference Pair Construction, (3) DPO Alignment. Overall, AGFSync learns from AI generated feedback data with DPO. AGFSync requires no human annotation, model architecture changes, or reinforcement learning.
The overall pipeline of our proposed AGFSync is illustrated as in Fig. 2.
@@ 3.1 Preference Candidate Set Generation
To encourage the diffusion model generate diverse style images for further text-image pair preference dataset, we intuitively think about using LLM to construct prompts that serve as image captions $\bm{c}$.
We employ LLM to generate prompts $\bm{c}$ from the instruction that would further feed into T2I diffusion model. We encourage the LLM to generate 12 distinct categories for diversity: Natural Landscapes, Cities and Architecture, People, Animals, Plants, Food and Beverages, Sports and Fitness, Art and Culture, Technology and Industry, Everyday Objects, Transportation, and Abstract and Conceptual Art. The established 45.8K dataset’s statistics is in Tab. 3.
@@
To construct the preference candidate set, we consider a text-conditioned generative diffusion model $G$ for candidate the image generation, where $G$ accept input parameters: text condition $\bm{c}$ and latent space noise $\bm{z}_{0}$. We let the diffusion model to generate $N_{\bm{c}}$ candidate images. To enhance the diversity and distinctiveness of the images produced by the model, we incorporate unit Gaussian noise into the conditional input $\bm{c}$ and generate $\bm{z}_{0}$ with different random seeds. This approach aims to introduce more randomness and variation to avoid overly uniform or similar generated images. Specifically, the process of generating backup images can be represented as in Eq. 1:
@@
Therefore, currently we have preference candidate $\bm{c}$ and its corresponding $N_{\bm{c}}$ generated images. We will select and refine these candidates for constructing the final preference pair dataset.
@@ 3.2.1 VQA Questions Generation
To make the score easier to calculate, we ensure that the answers to these questions are uniformly "yes" in the instruction prompt. To refine the questions, we let the LLM to validate the questions if they are ambiguous or unrelated to the captions, therefore all questions are generated not valid or closely related to the text for answering by the validation process in the instruction prompt.
3.2.2 VQA Score
@@
The VQA model $\Phi$ is employed to answer all questions ${Q_{i}(\bm{c})}_{i=1}^{N_{\bm{c}}}$ based on the image $\bm{x}_{0}$. The correctness of the VQA model’s answers is evaluated by comparing them to the correct answers $A_{i}(\bm{c})$. The VQA score [9], which quantifies the consistency between the text and the generated image, is calculated in Eq. 2:
@@
3.2.3 CLIP Score
Utilizing the CLIP [19] model, we convert the prompt words and the generated image into vector representations, denoted as $\bm{c}^{(emb)}$ for text and $\bm{x}^{\prime}_{0}$ for the image. The cosine similarity between the two vectors, computed in a shared embedding space, quantifies the alignment between the text and the image, embodying the CLIP Score [8]. The formula is refined as Eq. 3.
@@ 3.2.3 CLIP Score
3.2.4 Aesthetic Score
The aesthetic score assesses an image’s visual appeal by analyzing multifaceted elements like composition, color harmony, style, and high-level semantics, which collectively contribute to the aesthetic quality of an image [10]. The evaluation is represented in Eq. 4:
@@ 3.2.4 Aesthetic Score
3.2.5 Weighted Score Calculation
@@
3.2.6 Preference Pair Dataset Construction
With the previous generated set of $N$ images $\bm{X}_{0}=\{\bm{x}^{1}_{0},\bm{x}^{2}_{0},\dots,\bm{x}^{N}_{0}\}$ for a given textual prompt $\bm{c}$, each candidate image is then evaluated to assign the score calculated in multiple aspects as the aformentioned weighted score. To identify the most and least preferred images, which we term as the "winner" and "loser", we employ the following selection criteria in Eq. 6 and Eq. 7:
@@ 3.2.6 Preference Pair Dataset Construction
This approach yields a preference pair for each textual prompt $\bm{c}$, represented as $(\bm{c},\bm{x}^{w}_{0},\bm{x}^{l}_{0})$. The rationale behind selecting the highest and lowest scored images is to capture the widest possible discrepancy in quality and relevance, providing a clear contrast suitable for fine-tuning with DPO.
@@ 3.3 DPO Alignment
Derive from Diffusion-DPO [20], we consider the preference dataset, denoted as $\mathcal{D}=\{(\bm{c},\bm{x}^{w}_{0},\bm{x}^{l}_{0})\}$. Applying DPO for diffusion models is modeled as the following objective function Eq. 8. For the detailed notation of algorithms Eq. 8, please refer to Diffusion-DPO [20] and DPO [20].
$L_{\mathrm{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})}$ (8)
$\log\sigma(-\beta T\omega(\lambda_{t})(\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\theta}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}-\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}$
$-(\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\theta}(\bm{x}^{l}_{t},t,\bm{c})\|_{2}^{2}-\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{l}_{t},t,\bm{c})\|_{2}^{2})))$
where $\bm{x}^{*}_{t}=\alpha_{t}\bm{x}^{*}_{0}+\sigma_{t}\bm{\epsilon}^{*}$, $\bm{\epsilon}^{*}\sim\mathcal{N}(0,\bm{I})$. Here, $\alpha_{t}$ and $\sigma_{t}$ are the noise scheduling functions as defined in [23]. Consequently, $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$. Similar to [27], we incorporate $T$ and $\omega(\lambda_{t})$ into the constant $\beta$.

### v6

# AGFSync: Leveraging AI-Generated Feedback for Preference Optimization in Text-to-Image Generation

@@ Abstract
Text-to-Image (T2I) diffusion models have achieved remarkable success in image generation. Despite their progress, challenges remain in both prompt-following ability, image quality and lack of high-quality datasets, which are essential for refining these models. As acquiring labeled data is costly, we introduce AGFSync, a framework that enhances T2I diffusion models through Direct Preference Optimization (DPO) in a fully AI-driven approach. AGFSync utilizes Vision-Language Models (VLM) to assess image quality across style, coherence, and aesthetics, generating feedback data within an AI-driven loop. By applying AGFSync to leading T2I models such as SD v1.4, v1.5, and SDXL-base, our extensive experiments on the TIFA dataset demonstrate notable improvements in VQA scores, aesthetic evaluations, and performance on the HPSv2 benchmark, consistently outperforming the base models. AGFSync’s method of refining T2I diffusion models paves the way for scalable alignment techniques. Our code and dataset are publicly available.
Projecthttps://anjingkun.github.io/AGFSync
@@ 1 Introduction
The advent of Text-to-Image (T2I) generation technology represents a significant advancement in generative AI. Recent breakthroughs have predominantly utilized diffusion models to generate images from textual prompts (Rombach et al. 2022a; Betker et al. 2023; Podell et al. 2023; Zhang, Rao, and Agrawala 2023; Zhou et al. 2024b). However, achieving high fidelity and aesthetics in generated images poses challenges, including deviations from prompts and inadequate image quality (Zhang et al. 2023). Addressing these challenges requires enhancing diffusion models’ ability to accurately interpret detailed prompts (prompt-following ability (Betker et al. 2023)) and improve the generative quality across style, coherence, and aesthetics.
Efforts to overcome these challenges span dataset, model, and training levels. High-quality text-image pair datasets, as proposed in the data-centric AI philosophy, can significantly improve performance (Zhou et al. 2024a). Therefore a high-quality image caption and its corresponding image pair dataset is crucial in training (Betker et al. 2023).
@@
At the model architecture level, advancements include the optimization of cross-attention mechanisms to improve model compliance (Feng et al. 2023). These efforts, both at the dataset and model architecture levels, follow the traditional training paradigm of using elaborately designed models with specific datasets. In contrast, in the training domain, strategies inspired by the success of large language models, such as OpenAI’s ChatGPT (OpenAI 2023), include supervised finetuning (SFT) and alignment stages. With a pretrained T2I diffusion model, enhancing the model for better image quality can be approached in either the SFT stage or the alignment stage. The former approach, as seen in the latest work DreamSync (Sun et al. 2023), finetunes the diffusion model through a selected image selection procedure where a Vision-Language Model (VLM) (Achiam et al. 2023; Qin et al. 2023; Zhou et al. 2024c; Qin et al. 2024) evaluates and then selects high-quality text-image pairings for further finetuning. However, DreamSync exhibits a lower prompt generation conversion rate and is limited by the intrinsic capabilities of the diffusion model, leading to uncontrollable data distribution in the finetuning dataset. The latter approach, DPOK (Fan et al. 2024), DDPO (Black et al. 2023), and DPO (Rafailov et al. 2023) use reinforcement learning for alignment, while Diffusion-DPO (Wallace et al. 2023) applies Direct Preference Optimization (DPO) for model alignment, modifying the original DPO algorithm to directly optimize diffusion models based on preference data. Yet, it only focuses on evaluating image quality from one aspect. Furthermore, existing methods mostly depend on extensive, quality-controlled labeled data.
@@
Addressing these requires a cost-effective, low-labor approach that minimizes the need for human-labeled data while considering multiple quality aspects of images. Leveraging AI in generating datasets and evaluating image quality can fill these gaps without human intervention. Through generating diverse textual prompts, assessing generated images, and constructing a comprehensive preference dataset, AGFSync epitomizes the full spectrum of AI-driven innovation—ushering in an era of enhanced data utility, accessibility, scalability, and process automation while simultaneously mitigating the costs and limitations associated with manual data labeling.
@@
More specifically, AGFSync aligns text-to-image diffusion models via DPO, with multi-aspect AI feedback generated data. The process begins with the preference candidate set generation, where LLM generates descriptions of diverse styles and categories, serving as high-quality textual prompts. Candidate images are then generated using these AI-generated prompts, therefore constructing candidate prompt-image pairs. Image evaluation and VQA data construction follow, using LLM to generate questions related to the composition elements, style, etc., based on its initial prompts. VQA scoring is conducted by inputting these questions into the VQA model to assess whether the diffusion model-generated images aesthetically follow the prompts, calculating accuracy as the VQA score. With combined weighted scores of VQA, CLIP, and aesthetics filtering, the preference pair dataset is established within the best and worst images. Finally, DPO alignment is applied to the diffusion model using the constructed preference pair dataset. The entire process leverages the robust capabilities of VLMs without any human engagement, ensuring a human-free, cost-effective workflow.
@@
    Our proposed framework AGFSync, aided by multiple evaluation scores, leveraging DPO finetuning approach, introduces a fully automated, AI-driven approach, which elevates fidelity and aesthetic quality across varied scenarios without human annotations.
@@
    Extensive experiments demonstrate that AGFSync significantly and consistently improves upon existing diffusion models in terms of adherence to text prompts and overall image quality, establishing the efficacy and transformative potential of our AI-driven data generation, evaluation and finetuning framework.
@@ 2.1 Aligning Diffusion Models Methods
Dataset-Level Approaches: A pivotal aspect of improving image generation models involves curating and finetuning datasets that are deemed visually appealing. Works by (Podell et al. 2023; Rombach et al. 2022a) utilize datasets rated highly by aesthetics classifiers (Schuhmann 2022) to bias models towards generating visually appealing images. Similarly, Emu (Dai et al. 2023b) enhances both visual appeal and text alignment through finetuning on a curated dataset of high-quality photographs with detailed captions. Efforts to re-caption web-scraped image datasets for better text fidelity are evident in (Betker et al. 2023; Segalis et al. 2023). Moreover, similar to finetuning LLMs with generated data (Betker et al. 2023; Segalis et al. 2023), DreamSync (Sun et al. 2023) improves T2I synthesis with feedback from vision-language image understanding models, aligning images with textual input and the aesthetic quality of the generated images.
Model-Level Enhancements: At the model level, enhancing the architecture with additional components like attention modules (Feng et al. 2023) offers a training-free solution to enhance model compliance with desired outputs. StructureDiffusion (Feng et al. 2023) and SynGen (Rassin et al. 2023) also work on training-free methods that focus on model’s inference time adjustments.
Training-Level Strategies: The integration of supervised finetuning (SFT) with advanced alignment stages, such as reinforcement learning approaches like DPOK (Fan et al. 2024), DDPO (Black et al. 2023), and DPO (Rafailov et al. 2023), shows significant potential in aligning image quality with human preferences. Among these, Diffusion-DPO emerges as an RL-free method, distinct from other RL-based alignment strategies, effectively enhancing human appeal while ensuring distributional integrity (Wallace et al. 2023).
A common drawback of these approaches is the expensive finetuning dataset, as most of them rely on human-annotated data and human evaluation. This paradigm cannot support training an extensive and scalable diffusion model.
@@ 2.2 Image Quality Evaluation Methods
Evaluating image quality in a comprehensive manner is pivotal, integrating both automated benchmarks and human assessments to ensure fidelity and aesthetic appeal. The introduction of TIFA (Hu et al. 2023) utilize Visual Question Answering (VQA) models to measure the faithfulness of generated images to text prompts, setting a foundation for subsequent innovations. The CLIP score (Hessel et al. 2021) builds upon CLIP (Radford et al. 2021) enables a reference-free evaluation of image-caption compatibility through the computation of cosine similarity between image and text embeddings, showcasing high correlation with human judgments without needing reference captions. PickScore (Kirstain et al. 2024) leverages user preferences to predict the appeal of generated images, combining CLIP model elements with InstructGPT’s reward model objectives (Ouyang et al. 2022) for a nuanced understanding of user satisfaction. Alongside, the aesthetic score (Ke et al. 2023) assesses images based on aesthetics learned from image-comment pairs, providing a richer evaluation that includes composition, color, and style.
@@ 3 Methodology
Figure 1: Overall pipeline of AGFSync, which mainly encompasses 3 steps. AGFSync learns from AI-generated feedback data with DPO. AGFSync requires no human annotation, model architecture changes, or reinforcement learning.
The overall pipeline of AGFSync is illustrated in Fig. 1.
@@ 3.1 Preference Candidate Set Generation
To encourage the diffusion model to generate diverse style images for further text-image pair preference datasets, we employ LLM to construct prompts that serve as image captions $\bm{c}$.
We employ LLM to generate image captions $\bm{c}$ from the instruction that would further feed into the T2I diffusion model. We encourage the LLM to generate 12 distinct categories for diversity: Natural Landscapes, Cities and Architecture, People, Animals, Plants, Food and Beverages, Sports and Fitness, Art and Culture, Technology and Industry, Everyday Objects, Transportation, and Abstract and Conceptual Art.
@@
To construct the preference candidate set, we consider a text-conditioned generative diffusion model $G$ for candidate the image generation, where $G$ accept input parameters: text condition $\bm{c}$ and latent space noise $\bm{z}_{0}$. We let the diffusion model to generate $N$ candidate images. To enhance the diversity and distinctiveness of the images produced by the model, we incorporate Gaussian noise into the conditional input $\bm{c}$ and generate $\bm{z}_{0}$ with different random seeds. This approach aims to introduce more randomness and variation to avoid overly uniform or similar generated images. Specifically, the process of generating backup images can be represented as in Eq. 1:
@@
Therefore, we currently have the sample $\bm{c}$ and its corresponding $N$ preference candidate generated images. Next, we will filter and refine these candidates to construct the final preference pair dataset.
@@ 3.2 Preference Pair Construction
To make the score easier to calculate, we ensure that the answers to these questions are uniformly “yes” in the instruction prompt. To refine the questions, we let the LLM to validate the questions if they are ambiguous or unrelated to the captions, therefore all questions are generated not valid or closely related to the text for answering by the validation process in the instruction prompt.
VQA Score
@@
The VQA model $\Phi$ is employed to answer all questions ${Q_{i}(\bm{c})}_{i=1}^{N_{\bm{c}}}$ based on the image $\bm{x}_{0}$. The correctness of the VQA model’s answers is evaluated by comparing them to the correct answers $A_{i}(\bm{c})$. The VQA score (Hu et al. 2023), which quantifies the consistency between the text and the generated image, is calculated in Eq. 2:
@@
CLIP Score
Utilizing the CLIP (Radford et al. 2021) model, we convert the prompt words and the generated image into vector representations, denoted as $\bm{c}^{(emb)}$ for text and $\bm{x}^{\prime}_{0}$ for the image. The cosine similarity between the two vectors, computed in a shared embedding space, quantifies the alignment between the text and the image, embodying the CLIP Score (Hessel et al. 2021), defined in Eq. 3:
@@
Aesthetic Score
The aesthetic score assesses an image’s visual appeal by analyzing multifaceted elements like composition, color harmony, style, and high-level semantics, which collectively contribute to the aesthetic quality of an image (Ke et al. 2023). The evaluation is defined in Eq. 4:
@@
Weighted Score Calculation
@@
Preference Pair Dataset Construction
With the generated set of $N$ images $\bm{X}_{0}=\{\bm{x}^{1}_{0},\bm{x}^{2}_{0},\dots,\bm{x}^{N}_{0}\}$ for a given textual prompt $\bm{c}$, each candidate image is then evaluated to assign the score calculated in multiple aspects as the aformentioned weighted score. To identify the most and least preferred images, which termed as the “winner” and “loser”, we apply the selection criteria in Eq. 6 and Eq. 7:
@@
This approach yields a preference pair for each textual prompt $\bm{c}$, represented as $(\bm{c},\bm{x}^{w}_{0},\bm{x}^{l}_{0})$. The rationale behind selecting the highest and lowest scored images is to capture the widest possible discrepancy in quality and relevance, providing a clear contrast suitable for finetuning with DPO.
@@ 3.3 DPO Alignment
Derive from Diffusion-DPO (Rafailov et al. 2023), we consider the preference dataset, denoted as $\mathcal{D}=\{(\bm{c},\bm{x}^{w}_{0},\bm{x}^{l}_{0})\}$. Applying DPO for diffusion models is modeled as the following objective function $L(\theta)$ in Eq. 8. For the detailed notation of algorithms Eq. 8, please refer to Diffusion-DPO (Rafailov et al. 2023) and DPO (Rafailov et al. 2023).
$L(\theta)=-\mathbb{E}_{(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{D},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})}$ (8)
$\quad\quad\quad\log\sigma(-\beta T\omega(\lambda_{t})(\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\theta}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}$
$\quad\quad\quad-\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{w}_{t},t,\bm{c})\|_{2}^{2}-(\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\theta}(\bm{x}^{l}_{t},t,\bm{c})\|_{2}^{2}$
$\quad\quad\quad-\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\text{ref}}(\bm{x}^{l}_{t},t,\bm{c})\|_{2}^{2})))$
where $\bm{x}^{*}_{t}=\alpha_{t}\bm{x}^{*}_{0}+\sigma_{t}\bm{\epsilon}^{*}$, $\bm{\epsilon}^{*}\sim\mathcal{N}(0,\bm{I})$. Here, $\alpha_{t}$ and $\sigma_{t}$ are the noise scheduling functions as defined in (Rombach et al. 2022a). Consequently, $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$. Similar to (Wallace et al. 2023), we incorporate $T$ and $\omega(\lambda_{t})$ into the constant $\beta$.